# DenseVLM: A Retrieval and Decoupled Alignment Framework for Open-Vocabulary Dense Prediction

### v3

@@ Abstract
Pre-trained vision-language models (VLMs), such as CLIP, have demonstrated impressive zero-shot recognition capability, but still underperform in dense prediction tasks. Self-distillation recently is emerging as a promising approach for fine-tuning VLMs to better adapt to local regions without requiring extensive annotations. However, previous state-of-the-art approaches often suffer from significant ‘foreground bias’, where models tend to wrongly identify background regions as foreground objects. To alleviate this issue, we propose DenseVLM, a framework designed to learn unbiased region-language alignment from powerful pre-trained VLM representations. DenseVLM leverages the pre-trained VLM to retrieve categories for unlabeled regions and then decouples the interference between foreground and background features. This separation ensures accurate region-category alignment while maintaining semantic distinctions during training. We show that DenseVLM can directly replace the original VLM in open-vocabulary object detection and image segmentation methods, leading to notable performance improvements. Furthermore, it exhibits promising zero-shot scalability when training on more extensive and diverse datasets. Our code is available at https://github.com/HVision-NKU/DenseVLM.
@@ 1 Introduction
Open-vocabulary dense prediction, primarily including object detection [68, 15, 48, 29, 22] and image segmentation [13, 63, 38, 64, 5, 35], aims to identify regions or dense visual concepts of arbitrary categories as described by the text. Benefiting from the powerful pre-trained Vision-Language Models (VLMs), recent open-vocabulary approaches [74, 27, 64] for dense prediction have achieved great progress.
@@
Figure 1: Illustration of foreground bias. Previous methods [55, 71, 59] often produce similar foreground predictions for background regions, our approach effectively alleviates this issue.
Popular VLMs, such as CLIP [52] and EVA-CLIP [55], have exhibited remarkable zero-shot recognition abilities for global image understanding [19, 20, 37, 36, 12]. However, these models expose notable limitations in the understanding of local visual semantics, particularly in localizing and identifying small objects and background stuff [71, 74]. This limitation arises from the training manner of VLMs that align images with global text while neglecting the correspondences between local image regions and their specific text descriptions. To alleviate this issue, some studies use region-text or pseudo region-text pairs [71, 34, 70, 41] but these methods are limited by the high annotation cost and lack scalability for open-world scenes. In contrast, self-supervised approaches, such as CLIPSelf [59] and MaskEmbed [6], align region semantics using cropped image representations or reconstruct masked patch embeddings, respectively. These self-distillation approaches, which do not rely on annotated data, offer flexibility and scalability across a variety of datasets.
@@
Figure 2: Comparison of different VLMs. Unlike existing methods using (a) image-text contrastive learning [52], (b) region-text contrastive learning [71] or (c) self-distillation [59], our method leverages powerful model representations for region-language alignment.
@@
Despite the advances, previous VLMs [52, 55, 71] pre-trained on image-text and region-text pairs tend to overemphasize the foreground objects at the expense of the background context. This disproportionate focus results in a pronounced predilection for foreground object recognition and causes models to incorrectly associate background regions with foreground labels. Consequently, in dense prediction tasks, these VLMs often misclassify background regions to co-occurring foreground classes—a phenomenon we term as ‘foreground bias.’ To illustrate this issue, we present an example comparing the classification results of popular VLMs [55, 71, 59], where features are extracted from regions using ground-truth masks. As illustrated in Fig. 1, these models tend to confuse ‘sky’ (a background class) with ‘building’ (a foreground object).
@@
To address this bias, we propose aligning foreground and background regions separately, ensuring explicit semantic separation through distinct category sets. To achieve this, we introduce DenseVLM, an end-to-end framework designed for unbiased region-language alignment, as shown in Fig. 2(d). Specifically, for unlabeled regions, we leverage a powerful pre-trained VLM that has learned robust semantic features from diverse data to retrieve relevant categories without relying on paired data or self-distillation [55, 71, 59]. To ensure semantic diversity, DenseVLM incorporates a broad and comprehensive set of categories derived from large-scale datasets or predicted by generative models [67]. A key feature of DenseVLM is its ability to classify regions as either foreground or background based on predefined sets for each category. This classification way enables the decoupling of region features, reducing interactions between foreground and background. By decoupling these features, DenseVLM achieves independent region alignment of their respective categories while maintaining a semantic separation during training. Furthermore, DenseVLM improves both efficiency and performance by directly extracting region features from the dense features of VLMs, avoiding traditional image cropping [75, 59].
@@
We evaluate the effectiveness of DenseVLM on several open-vocabulary benchmarks [59], including object detection (box classification) and image segmentation (thing and stuff mask recognition). DenseVLM is adaptable to various network architectures, such as ViTs [9] and CNNs [32], and consistently outperforms other competing methods [52, 55, 59]. Furthermore, DenseVLM has great scaling ability, showing promising performance improvement when scaling up the training set based on the SA-1B [28] dataset. For downstream tasks, DenseVLM improves the two-stage models [29] of OV-COCO [3] and OV-LVIS [17] in open-vocabulary object detection and achieves notable gains in open-vocabulary semantic segmentation over state-of-the-art methods such as SAN [64] and Cat-Seg [5]. We summarize our contributions as follows.
@@ 2 Related work
Open-vocabulary dense prediction. Open-vocabulary dense prediction approaches aim to overcome the constraints of predefined categories, thereby enhancing their application in object detection [68, 15, 48, 29] and image segmentation [13, 63, 38, 64, 35, 33]. The success of pre-trained vision-language models like CLIP [52], has further inspired advancements in this area. In open-vocabulary detection, recent studies [15, 58] exploit the CLIP models to effectively identify novel objects. Furthermore, several works [29, 60] construct object detectors that utilize frozen CLIP encoders, thereby reducing computational overhead while maintaining performance. For open-vocabulary segmentation, a common two-stage pipeline [8, 7, 63, 43] integrates a class-agnostic mask generator with a fixed CLIP encoder to achieve cross-modal alignment and mask classification. Recent methods also investigate the use of shared frozen CLIP with side adapter networks [65] or the adoption of single-stage frameworks [66]. However, due to its training on image-text pairs, CLIP lacks precise local vision-language alignment, which is essential for dense prediction tasks. Although recent studies fine-tune CLIP for pixel-level [5, 62] or mask-level [24] alignment, they are constrained by the scarcity of densely labeled data.
@@
Vision-language alignment at image and region levels. Pre-training visual-language models enable alignment between image and text representations [21, 46, 52, 26, 47, 2, 57]. By using image-text pairs [54] as training data, methods such as CLIP [52] and ALIGN [21] exhibit impressive zero-shot classification capabilities. To improve the alignment of vision-language of dense features, some training-free studies [56, 30, 31, 61] attempt to modify the output layers of CLIP. Recent methods [50, 53, 51] align visual patches with image-level text or learn local-global correspondence via self-distillation, both requiring extensive text-image pairs.
For precise local region alignment, researchers utilize annotations from visual grounding datasets [41] to train models on region-text alignment. For example, some methods such as GLIP [34, 70] and Grounding DINO [41] learn region-language grounding from region-text pairs or generates pseudo region-text pairs like RegionCLIP [71]. In the context of open-vocabulary detection and segmentation, some works achieve dense visual and text alignment by using mask attention [8, 74, 27, 64] or fine-tuning CLIP [62, 5, 23]. However, these methods are constrained by the high cost of annotation, making large-scale deployment challenging. To address the issue of annotated data scarcity, recent methods like CLIPSelf [59] use cropped images for semantic alignment, while MaskEmbed [6] leverages masked patch embeddings for feature reconstruction. Despite these advances, the efficacy of self-distillation methods is constrained by the suboptimal performance of the teacher model and is further compromised by foreground bias. To overcome these limitations, we leverage a powerful VLM to retrieve categories for unlabeled regions while decoupling foreground and background features through textual category guidance.
@@ 3 Method
Our aim is to develop a region-language alignment model that can effectively align local visual and semantic features, thereby enhancing the performance of VLMs in open-vocabulary dense prediction tasks. To achieve this, it is crucial to alleviate the foreground bias problem prevalent in previous VLMs [52, 55] that arises from training on image-text pairs. Moreover, our approach seeks to receive better performance beyond the constraints of self-distillation [59].
@@ 3.1 VLM’s representation
VLMs are typically designed to learn both global visual and textual representations within a shared semantic space. In ViT-based VLMs [4, 55], dense visual features are extracted through residual attention blocks. Following [74, 59], we derive the dense image feature map $\mathcal{F}$ by modifying the final residual attention block, i.e., removing the SoftMax operation and incorporating mapping layers. For a set of categories $\{c^{1},\dots,c^{D}\}$, where $D$ is the total number of categories, textual descriptions are generated using a prompt-template strategy [15], like “This is a photo of the c{c} in the scene.” These prompts are then encoded into text embeddings $\mathcal{T}=\{\mathcal{T}^{1},\ldots,\mathcal{T}^{D}\}$ by the text encoder.
@@ 3.2 DenseVLM framework
We propose DenseVLM, an end-to-end region-language alignment framework designed to mitigate foreground bias. DenseVLM achieves this by precisely aligning unlabeled regions with their corresponding categories. As shown in Fig. 3, the framework consists of two key components. First, it retrieves category semantics for region features using the P-VLMs (Powerful VLMs) with frozen weights. Second, it decouples these region-language alignments into foreground and background to train U-VLMs (Unbiased VLMs) without foreground bias. In particular, DenseVLM operates without requiring any annotations, relying on diverse category semantics from large-scale datasets [1, 73] or generating category sets from images using the NXTP [67].
Image patches to semantic regions. To achieve region alignment, the patch-level visual features from Sec. 3.1 need to be transformed into semantic region features. We adopt a strategy similar to [59], dividing the dense feature map into an $m\times n$ grid of patches. Unlike this approach, we refrain from directly cropping the input image, thereby enhancing both computational efficiency and representation effectiveness. In each training iteration, $m$ and $n$ are randomly selected from the set $\{2,\cdots,M\}$, where $M$ defaults to 6, allowing for varying patch grid sizes. The semantic region representation $\{\mathcal{F}^{1},\cdots,\mathcal{F}^{m\times n}\}$ is then extracted from dense feature map $\mathcal{F}$ via pooling (RoIAlign) [18]. This patch sampling strategy effectively captures region features of both foreground objects and background scenes. However, due to the foreground bias in VLMs [52, 55, 59], we observe that background regions are usually misclassified as foreground classes, despite containing only a small proportion of foreground-related patches.
@@
OpenCLIP [4] 49.8 74.3 51.9 72.2 29.2 54.9 28.4 54.1 29.6 53.4 37.9 66.6
DFN [11] 38.3 65.0 31.0 57.0 26.4 54.9 30.6 57.9 24.2 49.9 32.2 57.7
SigLIP [69] 39.9 61.4 40.4 60.1 30.3 56.4 25.9 49.2 27.3 47.6 34.5 57.3
EVA-CLIP [55] 44.3 68.7 44.7 66.0 26.2 51.9 33.0 57.6 33.9 56.2 36.2 62.3
RegionCLIP [71] 68.5 89.5 60.7 84.3 22.0 53.5 43.2 72.2 34.0 62.6 37.7 68.6
FineCLIP [25] 64.7 86.1 62.5 80.9 36.9 70.3 43.9 71.2 45.5 68.6 46.0 74.8
CLIPSelf [59] 69.1 88.2 66.7 83.0 41.7 75.2 48.1 77.7 47.5 74.2 53.7 82.8
@@ 4.1 Benchmarks
Experiment settings. To verify the effectiveness of the proposed DenseVLM, we perform experiments on dense prediction tasks using the COCO panoptic [40] val2017 split and ADE20K panoptic [72] val split. Following CLIPSelf [59], we evaluate box classification using pooled box features (Boxes) and mask classification with pooled mask features, distinguishing foreground objects (Masks-T) from background content (Masks-S). This process follows a similar procedure as illustrated in Fig. 1, where we use ground truth annotations to extract local features and evaluate the local classification. Results are reported in Top-1 and Top-5 mean accuracy across all experiments.

### v4

# Unbiased Region-Language Alignment for Open-Vocabulary Dense Prediction

@@ Abstract
Pre-trained vision-language models (VLMs), such as CLIP, have demonstrated impressive zero-shot recognition capability, but still underperform in dense prediction tasks. Self-distillation recently is emerging as a promising approach for fine-tuning VLMs to better adapt to local regions without requiring extensive annotations. However, previous state-of-the-art approaches often suffer from significant ‘foreground bias’, where models tend to wrongly identify background regions as foreground objects. To alleviate this issue, we propose DenseVLM, a framework designed to learn unbiased region-language alignment from powerful pre-trained VLM representations. DenseVLM leverages the pre-trained VLM to retrieve categories for unlabeled regions and then decouples the interference between foreground and background features. This separation ensures accurate region-category alignment while maintaining semantic distinctions during training. We show that DenseVLM can directly replace the original VLM in open-vocabulary object detection and image segmentation methods, leading to notable performance improvements. Furthermore, it exhibits promising zero-shot scalability when training on more extensive and diverse datasets. Our code is publicly available https://github.com/HVision-NKU/DenseVLM.
@@ 1 Introduction
Open-vocabulary dense prediction, primarily including object detection [Zareian_2021_CVPR, gu2021open, minderer2022simple, kuo2022f, jia2024mssd] and image segmentation [openseg, odise, ovseg, san, catseg, li2024cascade], aims to identify regions or dense visual concepts of arbitrary categories as described by the text. Benefiting from the powerful pre-trained Vision-Language Models (VLMs), recent open-vocabulary approaches [maskclip1, Kim_2023_CVPR, san] for dense prediction have achieved great progress.
@@
Figure 1: Illustration of foreground bias. Previous methods [sun2023eva, zhong2022regionclip, wu2023clipself] often produce similar foreground predictions for background regions, our approach effectively alleviates this issue.
Popular VLMs, such as CLIP [CLIP] and EVA-CLIP [sun2023eva], have exhibited remarkable zero-shot recognition abilities for global image understanding [huang2024class, huang2025mindgappreservingcompensating, li2024advancing, li2024promptkd, fang2025aligning]. However, these models expose notable limitations in the understanding of local visual semantics, particularly in localizing and identifying small objects and background stuff [zhong2022regionclip, maskclip1]. This limitation arises from the training manner of VLMs that align images with global text while neglecting the correspondences between local image regions and their specific text descriptions. To alleviate this issue, some studies use region-text or pseudo region-text pairs [zhong2022regionclip, li2022grounded, zhang2022glipv2, liu2023grounding] but these methods are limited by the high annotation cost and lack scalability for open-world scenes. In contrast, self-supervised approaches, such as CLIPSelf [wu2023clipself] and MaskEmbed [covert2024locality], align region semantics using cropped image representations or reconstruct masked patch embeddings, respectively. These self-distillation approaches, which do not rely on annotated data, offer flexibility and scalability across a variety of datasets.
@@
Figure 2: Comparison of different VLMs. Unlike existing methods using (a) image-text contrastive learning [CLIP], (b) region-text contrastive learning [zhong2022regionclip] or (c) self-distillation [wu2023clipself], our method leverages powerful model representations for region-language alignment.
@@
Despite the advances, previous VLMs [CLIP, sun2023eva, zhong2022regionclip] pre-trained on image-text and region-text pairs tend to overemphasize the foreground objects at the expense of the background context. This disproportionate focus results in a pronounced predilection for foreground object recognition and causes models to incorrectly associate background regions with foreground labels. Consequently, in dense prediction tasks, these VLMs often misclassify background regions to co-occurring foreground classes—a phenomenon we term as ‘foreground bias.’ To illustrate this issue, we present an example comparing the classification results of popular VLMs [sun2023eva, zhong2022regionclip, wu2023clipself], where features are extracted from regions using ground-truth masks. As illustrated in Fig. 1, these models tend to confuse ‘sky’ (a background class) with ‘building’ (a foreground object).
@@
To address this bias, we propose aligning foreground and background regions separately, ensuring explicit semantic separation through distinct category sets. To achieve this, we introduce DenseVLM, an end-to-end framework designed for unbiased region-language alignment, as shown in Fig. 2(d). Specifically, for unlabeled regions, we leverage a powerful pre-trained VLM that has learned robust semantic features from diverse data to retrieve relevant categories without relying on paired data or self-distillation [sun2023eva, zhong2022regionclip, wu2023clipself]. To ensure semantic diversity, DenseVLM incorporates a broad and comprehensive set of categories derived from large-scale datasets or predicted by generative models [Yue_2024_CVPR]. A key feature of DenseVLM is its ability to classify regions as either foreground or background based on predefined sets for each category. This classification way enables the decoupling of region features, reducing interactions between foreground and background. By decoupling these features, DenseVLM achieves independent region alignment of their respective categories while maintaining a semantic separation during training. Furthermore, DenseVLM improves both efficiency and performance by directly extracting region features from the dense features of VLMs, avoiding traditional image cropping [zhu2024mrovseg, wu2023clipself].
@@
We evaluate the effectiveness of DenseVLM on several open-vocabulary benchmarks [wu2023clipself], including object detection (box classification) and image segmentation (thing and stuff mask recognition). DenseVLM is adaptable to various network architectures, such as ViTs [dosovitskiy2020image] and CNNs [lecun1989backpropagation], and consistently outperforms other competing methods [CLIP, sun2023eva, wu2023clipself]. Furthermore, DenseVLM has great scaling ability, showing promising performance improvement when scaling up the training set based on the SA-1B [kirillov2023segment] dataset. For downstream tasks, DenseVLM improves the two-stage models [kuo2022f] of OV-COCO [chen2015microsoft] and OV-LVIS [Gupta_2019_CVPR] in open-vocabulary object detection and achieves notable gains in open-vocabulary semantic segmentation over state-of-the-art methods such as SAN [san] and Cat-Seg [catseg]. We summarize our contributions as follows.
@@ 2 Related work
Open-vocabulary dense prediction. Open-vocabulary dense prediction approaches aim to overcome the constraints of predefined categories, thereby enhancing their application in object detection [Zareian_2021_CVPR, gu2021open, minderer2022simple, kuo2022f] and image segmentation [openseg, odise, ovseg, san, li2024cascade, li2024clip]. The success of pre-trained vision-language models like CLIP [CLIP], has further inspired advancements in this area. In open-vocabulary detection, recent studies [gu2021open, Wu_2023_CVPR] exploit the CLIP models to effectively identify novel objects. Furthermore, several works [kuo2022f, CORA] construct object detectors that utilize frozen CLIP encoders, thereby reducing computational overhead while maintaining performance. For open-vocabulary segmentation, a common two-stage pipeline [maskclip, zegformer, odise, liu2023open] integrates a class-agnostic mask generator with a fixed CLIP encoder to achieve cross-modal alignment and mask classification. Recent methods also investigate the use of shared frozen CLIP with side adapter networks [xu2023san] or the adoption of single-stage frameworks [fcclip]. However, due to its training on image-text pairs, CLIP lacks precise local vision-language alignment, which is essential for dense prediction tasks. Although recent studies fine-tune CLIP for pixel-level [catseg, xie2023sed] or mask-level [jiao2024collaborative] alignment, they are constrained by the scarcity of densely labeled data.
@@
Vision-language alignment at image and region levels. Pre-training visual-language models enable alignment between image and text representations [align, Vilbert, CLIP, jose2025dinov2, ma2024clip, chen2023vlp, wang2023large]. By using image-text pairs [schuhmann2022laion] as training data, methods such as CLIP [CLIP] and ALIGN [align] exhibit impressive zero-shot classification capabilities. To improve the alignment of vision-language of dense features, some training-free studies [wang2025sclip, lan2024clearclip, lan2024proxyclip, wysoczanska2024clip] attempt to modify the output layers of CLIP. Recent methods [mukhoti2023open, ranasinghe2023perceptual, naeem2024silc] align visual patches with image-level text or learn local-global correspondence via self-distillation, both requiring extensive text-image pairs.
For precise local region alignment, researchers utilize annotations from visual grounding datasets [liu2023grounding] to train models on region-text alignment. For example, some methods such as GLIP [li2022grounded, zhang2022glipv2] and Grounding DINO [liu2023grounding] learn region-language grounding from region-text pairs or generates pseudo region-text pairs like RegionCLIP [zhong2022regionclip]. In the context of open-vocabulary detection and segmentation, some works achieve dense visual and text alignment by using mask attention [maskclip, maskclip1, Kim_2023_CVPR, san] or fine-tuning CLIP [xie2023sed, catseg, jiao2023learning]. However, these methods are constrained by the high cost of annotation, making large-scale deployment challenging. To address the issue of annotated data scarcity, recent methods like CLIPSelf [wu2023clipself] use cropped images for semantic alignment, while MaskEmbed [covert2024locality] leverages masked patch embeddings for feature reconstruction. Despite these advances, the efficacy of self-distillation methods is constrained by the suboptimal performance of the teacher model and is further compromised by foreground bias. To overcome these limitations, we leverage a powerful VLM to retrieve categories for unlabeled regions while decoupling foreground and background features through textual category guidance.
@@ 3 Method
Our aim is to develop a region-language alignment model that can effectively align local visual and semantic features, thereby enhancing the performance of VLMs in open-vocabulary dense prediction tasks. To achieve this, it is crucial to alleviate the foreground bias problem prevalent in previous VLMs [CLIP, sun2023eva] that arises from training on image-text pairs. Moreover, our approach seeks to receive better performance beyond the constraints of self-distillation [wu2023clipself].
@@ 3.1 VLM’s representation
VLMs are typically designed to learn both global visual and textual representations within a shared semantic space. In ViT-based VLMs [openclip, sun2023eva], dense visual features are extracted through residual attention blocks. Following [maskclip1, wu2023clipself], we derive the dense image feature map $\mathcal{F}$ by modifying the final residual attention block, i.e., removing the SoftMax operation and incorporating mapping layers. For a set of categories $\{c^{1},\dots,c^{D}\}$, where $D$ is the total number of categories, textual descriptions are generated using a prompt-template strategy [gu2021open], like “This is a photo of the c{c} in the scene.” These prompts are then encoded into text embeddings $\mathcal{T}=\{\mathcal{T}^{1},\ldots,\mathcal{T}^{D}\}$ by the text encoder.
@@ 3.2 DenseVLM framework
We propose DenseVLM, an end-to-end region-language alignment framework designed to mitigate foreground bias. DenseVLM achieves this by precisely aligning unlabeled regions with their corresponding categories. As shown in Fig. 3, the framework consists of two key components. First, it retrieves category semantics for region features using the P-VLMs (Powerful VLMs) with frozen weights. Second, it decouples these region-language alignments into foreground and background to train U-VLMs (Unbiased VLMs) without foreground bias. In particular, DenseVLM operates without requiring any annotations, relying on diverse category semantics from large-scale datasets [caesar2018coco, ade20k] or generating category sets from images using the NXTP [Yue_2024_CVPR].
Image patches to semantic regions. To achieve region alignment, the patch-level visual features from Sec. 3.1 need to be transformed into semantic region features. We adopt a strategy similar to [wu2023clipself], dividing the dense feature map into an $m\times n$ grid of patches. Unlike this approach, we refrain from directly cropping the input image, thereby enhancing both computational efficiency and representation effectiveness. In each training iteration, $m$ and $n$ are randomly selected from the set $\{2,\cdots,M\}$, where $M$ defaults to 6, allowing for varying patch grid sizes. The semantic region representation $\{\mathcal{F}^{1},\cdots,\mathcal{F}^{m\times n}\}$ is then extracted from dense feature map $\mathcal{F}$ via pooling (RoIAlign) [he2017mask]. This patch sampling strategy effectively captures region features of both foreground objects and background scenes. However, due to the foreground bias in VLMs [CLIP, sun2023eva, wu2023clipself], we observe that background regions are usually misclassified as foreground classes, despite containing only a small proportion of foreground-related patches.
@@
OpenCLIP [openclip] 49.8 74.3 51.9 72.2 29.2 54.9 28.4 54.1 29.6 53.4 37.9 66.6
DFN [fang2023data] 38.3 65.0 31.0 57.0 26.4 54.9 30.6 57.9 24.2 49.9 32.2 57.7
SigLIP [siglip] 39.9 61.4 40.4 60.1 30.3 56.4 25.9 49.2 27.3 47.6 34.5 57.3
EVA-CLIP [sun2023eva] 44.3 68.7 44.7 66.0 26.2 51.9 33.0 57.6 33.9 56.2 36.2 62.3
RegionCLIP [zhong2022regionclip] 68.5 89.5 60.7 84.3 22.0 53.5 43.2 72.2 34.0 62.6 37.7 68.6
FineCLIP [jing2025fineclip] 64.7 86.1 62.5 80.9 36.9 70.3 43.9 71.2 45.5 68.6 46.0 74.8
CLIPSelf [wu2023clipself] 69.1 88.2 66.7 83.0 41.7 75.2 48.1 77.7 47.5 74.2 53.7 82.8
@@ 4.1 Benchmarks
Experiment settings. To verify the effectiveness of the proposed DenseVLM, we perform experiments on dense prediction tasks using the COCO panoptic [lin2014microsoft] val2017 split and ADE20K panoptic [zhou2017scene] val split. Following CLIPSelf [wu2023clipself], we evaluate box classification using pooled box features (Boxes) and mask classification with pooled mask features, distinguishing foreground objects (Masks-T) from background content (Masks-S). This process follows a similar procedure as illustrated in Fig. 1, where we use ground truth annotations to extract local features and evaluate the local classification. Results are reported in Top-1 and Top-5 mean accuracy across all experiments.